Worked out solution

Problem 2/31 2/45 The cone falling with a speed vo strikes and pene- trates the block of packing material. The accelera- tion of the cone after impact is a g c2, where c is a positive constant and y is the penetration dis- tance. If the maximum penetration depth is ob- served to be ym, determine the constant c. Uo Problem 2/45
Answer #1

we know that acceleration is given by:

a = V*(dV/dy)

V*dV = a*dy

Now given that

a = g - cy^2

V_initial = initial velocity = v0

V_final = final velocity = 0 m/sec

max penetration depth = y_max = ym

And we need to find c = ?

So Using above equation:

V*dV = a*dy

Integrating both sides

(limit for velocity is v = v0 to v = 0 & limit for depth is y = 0 to y = ym)

a dy

int_{v_{0}}^{0}v*dv = int_{0}^{y_{m}}(g - cy^2)*dy

су

02 6 clym) c(0) 3

-20 = gym-

6

cy_{m}^3 = 3gy_{m} + rac{3v_{0}^2}{2}

2 5

300

c = rac{3}{2y_{m}^3}[2gy_{m} + v_{0}^2]

Please Upvote.

Comment below if you have any query.